A cubic sample of a new kind of artificial tissue is subject to an increase in pressure of 160 kPa which results in a reduction in the side length of the cube of 6.5%. That is, the side length of the cube has reduced from L0 to 0.935L0.

What is the bulk modulus of this tissue? (to 2 s.f and in MPa)
1
Expert's answer
2020-04-23T12:46:30-0400
"K=-\frac{\Delta P}{\frac{\Delta V}{V}}=\frac{\Delta P}{1-\frac{L^3}{L^3_0}}"

"K=\frac{0.16}{1-0.935^3}=0.88\ MPa"